# Towards a Theory of Pragmatic Information, with Applications

### v3

@@ Abstract
Standard information theory says nothing about how much meaning is conveyed by the information. We fill this gap with a rigorously justifiable, quantitative definition of “pragmatic information” as the amount of meaning in a message that is relevant to a particular decision (as opposed to the value that the decision maker places on that message). We posit that such a message updates a “state of the world” random variable, $\omega$, that informs the decision. The pragmatic information of a single message is then defined as the Kulbach-Leibler divergence between the a priori and updated probability distributions of $\omega$; the pragmatic information of a message ensemble is defined as the expected value of the pragmatic information values of the ensemble’s component messages. We then show, first, that the pragmatic information of a single message is the expected difference between the shortest binary encoding of $\omega$ under the a priori and updated distributions, and, second, that the average of the pragmatic values of individual messages, when sampled a large number of times from the ensemble, approaches its expected value.
Pragmatic information thus defined has many hoped-for properties, such as non-negativity and additivity for independent decisions and “pragmatically independent” messages. We also sketch three applications: The first is the single play of a slot machine, a.k.a. “one armed bandit”, with an unknown payout probability; the second being a characterization of the rate of biological evolution in a simple model of reproduction with mutation (the so-called “quasi-species” model); the third being a reformulation of the efficient market hypothesis of financial economics. We conclude by noting the importance of the computational capacity of the receiver in processing pragmatic information.
@@ 1 Introduction
Just about any discussion of standard information theory notes that the standard entropy measure of the amount of information in a transmitted message says nothing about the “meaningfulness” of the message. Indeed, Warren Weaver’s introduction to Shannon’s foundational paper [1] famously observes that the effectiveness of a communications process could be measured by answering any of the following three questions:
@@
Although the primary goal of the present work is to advance the theory as such, any theory that is termed “pragmatic” should have practical applications, so we mention a few. [2] proposed that the rate at which biological evolution aggregates “meaningful information” could be used to measure the rate at which evolution proceeds. Throughout the present paper, we present a second application, namely the role that the history of previous payouts of a slot machine acquire meaning in the context of deciding whether to keep playing. We also show how our conceptual framework makes possible a reformulation of the efficient market hypothesis of financial economics as the claim that none of the available information about a given security is “pragmatically useful”.
@@
We conclude with speculation about how a kind of “computational capacity” may play a role in the theory of pragmatic information similar to that of channel capacity in the Shannon theory, in that beliefs about $\omega$ cannot be updated during the non-zero time required to complete message processing. Given that a participant in the financial markets can only find market information useful if it is meaningful in our sense, we also discuss our ideas in the context of the celebrated efficient market hypothesis.
@@ 2.1 Our Conceptual Framework
While our conceptual framework can be applied to the wealth of so-called “multi-armed bandit problems” (See, for example, [9] and [10].), we consider the following, even simpler “one-armed bandit” problem: we can pay \$1 to play a slot machine (also known as a “one armed bandit”) as many times as we wish. After each play, we will win a \$2 payout with some constant probability $p$, or nothing, with probability $1-p$, with the payout events independent from one play to the next. The problem is that $p$ is not known in advance!
@@
Clearly, we learn something about $p$, and thus the correct play/no play decision for subsequent trials from the realized values of the sequence $\mathbf{1}_{1},\mathbf{1}_{2},\ldots,\mathbf{1}_{T}$ , but exactly how much? If we have won $w$ times in the $T$ trials, Laplace’s Rule of Succession [12] estimates $p$ as
$\hat{p}(w,T)=\frac{w+1}{T+2}.$ (1)
$\hat{p}(w,T)$ is less and less sensitive to whether or not any individual trial results in a payout as $T$ increases. Thus, we learn less and less about what we need to know from each additional value in this sequence; indeed, after a finite number of such values, we know essentially all we need to know to make our decision.
@@
$\mathcal{H}(\mathbf{1}_{t})=p\log_{2}p+(1-p)\log_{2}(1-p),$
per play remains constant from one play to the next, so that the Shannon information for the $T$ independent plays is $T[p\log_{2}p+(1-p)\log_{2}(1-p)]$. However, the Shannon information captures data such as the particular order of successes and failures, which is more than we need to know to estimate $p$.
@@ 2.2 Defining the Pragmatic Information of a Single Message
The pragmatic information, $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$, of a single message $m$ acting on $\Delta$ is [^3] As is customary in writing about information theory, all logarithms will subsequently be assumed to be base 2 logarithms.
@@
In our framework, $\Delta$ would be the shopkeeper who, presumably, needs to make decisions based on the true value of total sales for the day, which we take as the relevant random variable $\omega$. $\Delta$ would have an a priori estimate of $\omega$, based on the list of individual sales that is reflected in an a priori probability distribution, $\mathbf{q}$. This distribution, when updated upon learning the cash register output of total sales, $S_{T}$, for the day, is likely to become more peaked around a single value. Thus, if $\mathbf{p}_{T}$ is this updated distribution, $D_{\Delta}(\mathbf{p}_{T}||\mathbf{q})>0$.
@@ 2.3 Properties of the Pragmatic Information of a Single Message
The maximum value of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is $\max_{l}(-\log q_{l})$, where the max is taken over all of the $q$’s for which $p_{l|m}>0$.
@@ 3 The Pragmatic Information of an Ensemble of Messages
We now assume that the input messages are sampled from some finite ensemble $\mathcal{M}$ of possible messages, with message $m$ sampled with probability $\varphi_{m}$. We further assume that $\Delta$ makes a sequence of $N$ independent decisions [^5] We stress that, as in the previous section, $\Delta$ makes each of these decisions as though none of the previous decisions had ever been made., each with the same prior, and each informed by input message $M_{k}$ for $1\leq k\leq N$, with each $M_{k}$ drawn independently from $\mathcal{M}$ [^6] Here and subsequently, we use the symbol $M$ for a message sampled probabilistically from $\mathcal{M}$, so that $M$ is a random variable. In contrast, we use $m$ for deterministically identifying specific elements of $\mathcal{M}$, as we must in the summations below.. Given the additivity of Kullbach-Leibler divergence for independent decisions, the pragmatic information per message is the random variable
@@
Even though the terms in the above sum have different means, Kolmogorov’s Strong Law of Large Numbers [15] still applies because, per the corollary to Theorem 2.3.4, each term in this sum, and thus its variance, is bounded. We conclude that
@@
with probability 1. We therefore propose the following
@@
$p_{i|m}$ remains the conditional probablilty of outcome $\omega_{i}$, given the receipt of message $m$, so that the joint probability, $p_{i,m}$, of outcome $\omega_{i}$ upon receiving message $m$ is $p_{i,m}=\varphi_{m}p_{i|m}$. Once again, we take $p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)=0$ if $p_{i|m}=0$.
@@ 3.1 An Example: The One Armed Bandit Problem
It is instructive to apply this definition to the “one armed bandit problem” discussed above. Evidently, the relevant decision is whether it is worth playing the slot machine, a decision that is informed by a random variable, $\omega$, that can assume exactly one of the two values $\mathtt{PAYOUT}$ and $\mathtt{NOPAYOUT}$, i.e. $\Omega=\{\mathtt{PAYOUT},\mathtt{NOPAYOUT}\}$. We want to know the pragmatic information accruing from a knowledge of $\mathbf{1}_{T+1}$, given that the $T$ prior trials have resulted in $w$ wins. Thus, per (1), we take $\hat{p}(w,T)$ as $q_{1}$, our a priori estimate of the payout probability, and $1-\hat{p}(w,T)$ as $q_{0}$, our a priori estimate of the no payout probability.
@@
$D_{\Delta}(\mathbf{p}_{1}||\mathbf{q})$ $=\hat{p}(w+1,T+1)\log\left[\frac{\hat{p}(w+1,T+1)}{\hat{p}(w,T)}\right]+$
$\qquad\left[1-\hat{p}(w+1,T+1)\right]\log\left[\frac{1-\hat{p}(w+1,T+1)}{1-\hat{p}(w,T)}\right]$
@@
$D_{\Delta}(\mathbf{p}_{0}||\mathbf{q})$ $=\hat{p}(w,T+1)\log\left[\frac{\hat{p}(w,T+1)}{\hat{p}(w,T)}\right]+$
$\qquad\left[1-\hat{p}(w,T+1)\right]\log\left[\frac{1-\hat{p}(w,T+1)}{1-\hat{p}(w,T)}\right]$
@@
both of which can be interpreted as the reduction in the uncertainty surrounding the estimate of $p$, upon receipt of the corresponding value of $\mathbf{1}_{T+1}$. We conclude that the pragmatic information provided by the $T+1^{st}$ trial is
$\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ $=pD_{\Delta}(\mathbf{p}_{1}||\mathbf{q})+(1-p)D_{\Delta}(\mathbf{p}_{0}||\mathbf{q})$
@@
$\qquad+(1-p)\left\{\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right]\right\}$
As $T$ gets large, $w=pT+o(T)$ with probability 1, by the Kolmogorov Strong Law of Large Numbers. Straightforward algebraic manipulation then shows that the arguments of all of the logarithms in $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ approach 1 monotonically as $T$ increases. We conclude that $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ steadily decreases to zero with increasing $T$, as claimed above.
@@ 3.2 Properties of the Pragmatic Information of an Ensemble of Messages
$\Phi_{\Delta}(\mathcal{M};\Omega)\leq-\sum_{m}\varphi_{m}\log q_{k(m)},$
where, for each $m$, $k(m)=\operatorname*{arg\,max}(-\log q_{l})$, where the max is taken over all of the $q$’s for which $p_{l|m}>0$
@@
Next, we generalize Kullbach-Leibler Divergence. Suppose:
@@ 3.3 Pragmatic Independence of Messages
even though $M$ and $M^{\prime}$ are probabilistically independent.
3.4 The Relationship between Pragmatic Information and Mutual Information
This subsection clarifies the relationship between pragmatic information and the mutual information, $\mathcal{I}(\mathcal{M};\Omega)$, between $\mathcal{M}$ and $\Omega$. In particular, we resolve a confusion in [2] by showing that these two quantities are equal only in the important special case in which the a priori probabilities are equal to the marginal probabilties $p_{i}=\sum_{m\in\mathcal{M}}p_{i,m}$ for all $i$. Defining $\mathbf{p}$ as the vector of these marginal probabilities [^7] Note the distinction between $\mathbf{p}_{m}$ and $\mathbf{p}$: the former is the probability distribution of $\omega$, given a specific $m$, but the latter is the average of these distributions over all $m$’s. we have the following
@@ 3.4 The Relationship between Pragmatic Information and Mutual Information
$=\sum_{i,m}p_{i,m}\log\left[\frac{p_{i|m}}{p_{i}}\right]+\sum_{i}p_{i}\log\left[\frac{p_{i}}{q_{i}}\right]$
$=\mathcal{I}(\mathcal{M};\Omega)+D(\mathbf{p}||\mathbf{q}),$
@@
$\Phi(\mathcal{M};\Omega)\geq\mathcal{I}(\mathcal{M};\Omega),$ with equality if and only if $\mathbf{p}=\mathbf{q}$.

### v5

@@ Abstract
Standard information theory says nothing about how much meaning is conveyed by a message. We fill this gap with a rigorously justifiable, quantitative definition of “pragmatic information”, the amount of meaning in a message relevant to a particular decision. We posit that such a message updates a random variable, $\omega$, that informs the decision. The pragmatic information of a single message is then defined as the Kulbach-Leibler divergence between the apriori and aposteriori probabilities of $\omega$; the pragmatic information of a message ensemble is the expected value of the pragmatic information of the ensemble’s component messages. We justify these definitions by proving that the pragmatic information of a single message is the expected difference between the shortest binary encoding of $\omega$ under the a priori and a posteriori distributions, and that the average of the pragmatic values of individual messages, when sampled a large number of times from the ensemble, approaches its expected value.
Pragmatic information is non-negative and additive for independent decisions and “pragmatically independent” messages. Also, pragmatic information is the information analogue of free energy: just as free energy quantifies the part of a system’s total energy available to do useful work, so pragmatic information quantifies the information actually used in making a decision.
We sketch 3 applications: the single play of a slot machine, a.k.a. a “one armed bandit”, with an unknown payout probability; a characterization of the rate of biological evolution in the so-called “quasi-species” model; and a reformulation of the efficient market hypothesis of finance. We note the importance of the computational capacity of the receiver in each case.
@@ 1 Introduction
Just about any discussion of standard information theory notes that the standard entropy measure of the amount of information in a transmitted message says nothing about how meaningful the message is, even though a non-technical person would say that a message is “informative” if they are able to extract meaning from it. Warren Weaver’s introduction to Shannon’s foundational paper [2] addresses this point by observing that the effectiveness of a communications process could be measured by answering any of the following three questions:
@@
Although the primary goal of the present work is to advance the theory as such, any theory that is termed “pragmatic” should have practical applications, so we mention a few. [3] proposed that the rate at which biological evolution aggregates “meaningful information” could be used to measure the rate at which evolution proceeds. Throughout the present paper, we present a second application, namely the role that the history of previous payouts of a slot machine acquire meaning in the context of deciding whether to keep playing. We also show how our conceptual framework makes possible a reformulation of the celebrated efficient market hypothesis of financial economics as the claim that none of the available information about a given security is “pragmatically useful”.
@@
We conclude with speculation about how a kind of “computational capacity” may play a role in the theory of pragmatic information similar to that of channel capacity in the Shannon theory, in that beliefs about $\omega$ cannot be updated during the non-zero time required to complete message processing. Given that a participant in the financial markets can only find market information useful if it is meaningful in our sense, we also discuss our ideas in the context of the efficient market hypothesis.
@@ 2.1 Our Conceptual Framework
While our conceptual framework can be applied to the wealth of so-called “multi-armed bandit problems” (See, for example, [11] and [12].), we consider the following, even simpler “one-armed bandit” problem: we can pay \$1 to play a slot machine (also known as a “one armed bandit”) as many times as we wish. After each play, we will win a \$2 payout with some constant probability $\pi$, or nothing, with probability $1-\pi$, with the payout events independent from one play to the next. The problem is that $\pi$ is not known in advance!
@@
Clearly, we learn something about $\pi$, and thus the correct play/no play decision for subsequent trials from the realized values of the sequence $\mathbf{1}_{1},\mathbf{1}_{2},\ldots,\mathbf{1}_{T}$ , but exactly how much? If we have won $w$ times in the $T$ trials, Laplace’s Rule of Succession [14] estimates $\pi$ as
$\hat{\pi}(w,T)=\frac{w+1}{T+2}.$ (1)
$\hat{\pi}(w,T)$ is less and less sensitive to whether or not any individual trial results in a payout as $T$ increases. Thus, we learn less and less about what we need to know from each additional value in this sequence; indeed, after a finite number of such values, we know essentially all we need to know to make our decision.
@@
$\mathcal{H}(\mathbf{1}_{t})=\pi\log_{2}\pi+(1-p)\log_{2}(1-\pi),$
per play remains constant from one play to the next, so that the Shannon information for the $T$ independent plays is $T[\pi\log_{2}\pi+(1-\pi)\log_{2}(1-\pi)]$. However, the Shannon information captures data such as the particular order of successes and failures, which is more than we need to know to estimate $\pi$.
@@ 2.2 Defining the Pragmatic Information of a Single Message
The pragmatic information, $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$, of a single message $m$ acting on $\Delta$ is [^3] As is customary in writing about information theory, we assume base 2 logarithms here and subsequently.
@@
In our framework, $\Delta$ would be the shopkeeper who, presumably, needs to make decisions based on the true sales volume for the day, which we take as the relevant random variable $\omega$. $\Delta$ would have an a priori estimate of $\omega$, based on the list of individual sales that is reflected in an a priori probability distribution, $\mathbf{q}$. This distribution, when updated upon learning the cash register output of total sales, $S_{T}$, for the day, is likely to become more peaked around a single value. Thus, if $\mathbf{p}_{T}$ is this updated distribution, $D_{\Delta}(\mathbf{p}_{T}||\mathbf{q})>0$.
@@ 2.3 Properties of the Pragmatic Information of a Single Message
The maximum value of $D_{\Delta}(\mathbf{p}_{m}||\mathbf{q})$ is $\max_{l}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$.
@@ 3 The Pragmatic Information of an Ensemble of Messages
We now assume that $\Delta$ is asked to make a series of independent decisions, each time given the same prior set of beliefs (i.e. $\mathbf{q}$ is the same each time), but each time given a (possibly different) message $M_{k}$ from some finite ensemble $\mathcal{M}$. We further assume that each $M_{k}$ is sampled using the same probability distribution, namely ${\rm Pr}\left\{M_{k}=m\right\}=\varphi_{m}$ for all $k$ and every $m\in\mathcal{M}$ is eventually sampled, regardless of how the sampling starts. Given the additivity of Kullbach-Leibler divergence for independent decisions, $M_{k}$ conveys the (possibly different) amount of pragmatic information $D_{\Delta}(\mathbf{p}_{M_{k}}||\mathbf{q})$ for each $k$. The average amount of pragmatic information that $\Delta$ accrues per message is then the random variable
@@
Our assumptions on how the $M_{k}$’s are sampled are sufficient to guarantee that this sum satisfies Birkoff’s ergodic theorem [10]. We can therefore conclude that
@@
with probability 1, with the subscript being omitted from the expectation because the distribution of the $M_{k}$’s is the same for all $k$. We stress that, as in the previous section, $\Delta$ makes each of its decisions as though none of the previous decisions had ever been made. However, the hypotheses of the ergodic theorem are sufficiently weak that no such assumption is required for the $M_{k}$’s. While they could be sampled independently, they could also be a realization of a Markov chain, etc. The generality of this result therefore justifies the following
@@
where $p_{i|m}$ remains the conditional probablilty of outcome $\omega_{i}$, given the receipt of message $m$, so that the joint probability, $p_{i,m}$, of outcome $\omega_{i}$ upon receiving message $m$ is $p_{i,m}=\varphi_{m}p_{i|m}$. Once again, we take $p_{i|m}\log\left(\frac{p_{i|m}}{q_{i}}\right)=0$ if $p_{i|m}=0$.
@@ 3.1 An Example: The One Armed Bandit Problem
It is instructive to apply this definition to the “one armed bandit problem” discussed above. Evidently, the relevant decision is whether it is worth playing the slot machine, a decision that is informed by a random variable, $\omega$, that can assume exactly one of the two values $\mathtt{PAYOUT}$ and $\mathtt{NOPAYOUT}$, i.e. $\Omega=\{\mathtt{PAYOUT},\mathtt{NOPAYOUT}\}$. We want to know the pragmatic information accruing from a knowledge of $\mathbf{1}_{T+1}$, given that the $T$ prior trials have resulted in $w$ wins. Thus, per (1), we take $\hat{\pi}(w,T)$ as $q_{1}$, our a priori estimate of the payout probability, and $1-\hat{\pi}(w,T)$ as $q_{0}$, our a priori estimate of the no payout probability.
@@
$D_{\Delta}(\mathbf{\pi}_{1}||\mathbf{q})$ $=\hat{\pi}(w+1,T+1)\log\left[\frac{\hat{\pi}(w+1,T+1)}{\hat{\pi}(w,T)}\right]+$
$\qquad\left[1-\hat{\pi}(w+1,T+1)\right]\log\left[\frac{1-\hat{\pi}(w+1,T+1)}{1-\hat{\pi}(w,T)}\right]$
@@
$D_{\Delta}(\mathbf{\pi}_{0}||\mathbf{q})$ $=\hat{\pi}(w,T+1)\log\left[\frac{\hat{\pi}(w,T+1)}{\hat{\pi}(w,T)}\right]+$
$\qquad\left[1-\hat{\pi}(w,T+1)\right]\log\left[\frac{1-\hat{\pi}(w,T+1)}{1-\hat{\pi}(w,T)}\right]$
@@
both of which can be interpreted as the reduction in the uncertainty surrounding the estimate of $\pi$, upon receipt of the corresponding value of $\mathbf{1}_{T+1}$. We conclude that the pragmatic information provided by the $T+1^{st}$ trial is
$\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ $=\pi D_{\Delta}(\mathbf{\pi}_{1}||\mathbf{q})+(1-\pi)D_{\Delta}(\mathbf{\pi}_{0}||\mathbf{q})$
@@
$\qquad+(1-\pi)\left\{\frac{w+1}{T+3}\log\left[\frac{T+2}{T+3}\right]+\frac{T-w+2}{T+3}\log\left[\frac{(T-w+2)(T+2)}{(T+3)(T-w+1)}\right]\right\}$
As $T$ gets large, $w=\pi T+o(T)$ with probability 1, by the Kolmogorov Strong Law of Large Numbers [17]. Straightforward algebraic manipulation then shows that the arguments of all of the logarithms in $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ approach 1 monotonically as $T$ increases. We conclude that $\Phi_{\Delta}\left(\mathcal{M}_{T+1};\Omega\right)$ steadily decreases to zero with increasing $T$.
@@ 3.2 Properties of the Pragmatic Information of an Ensemble of Messages
$\Phi_{\Delta}(\mathcal{M};\Omega)$ $\leq-\sum_{m}\varphi_{m}\log q_{k(m)}$
$\leq-\log q_{l}$
where, for each $m$, $k(m)=\operatorname*{arg\,max}(-\log q_{l})$, with the max taken over all of the $q$’s for which $p_{l|m}>0$ and $-\log q_{l}$ is the largest of the $-\log q$’s (The $\varphi$’s sum to 1.).
@@
Next, we generalize joint and conditional Kullbach-Leibler Divergence. Suppose:
@@ 3.3 Pragmatic Independence of Messages
even though $M$ and $M^{\prime}$ of that subsection are probabilistically independent.
4 Pragmatic Information as “Free Information,” By Way of Mutual Information
Pragmatic information can be interpreted as a kind of information analog of the Hemholtz free energy of statistical mechanics in the following sense: the Hemholtz free energy, $F$, is given by the formula
$F=U-TS.$
It is is a measure of the amount of energy in a system available to do useful work at a constant temperature, $T$, as opposed to the total energy, $U$, of the system ($S$ is the system’s entropy). Similarly, as we now show, pragmatic information is what remains when information irrelevant to making the decision is removed.
Denote the mutual information between $\mathcal{M}$ and $\Omega$ as $\mathcal{I}(\mathcal{M};\Omega)$, $p_{i}=\sum_{m\in\mathcal{M}}p_{i,m}$ for all $i$, and $\mathbf{p}$ as the vector of these marginal probabilities [^5] Note the distinction between $\mathbf{p}_{m}$ and $\mathbf{p}$: the former is the probability distribution of $\omega$, given a specific $m$, but the latter is the average of these distributions over all $m$’s.. We then have the following
@@ 4 Pragmatic Information as “Free Information,” By Way of Mutual Information
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
@@
$=\sum_{i,m}p_{i,m}\log\left[\frac{p_{i|m}}{p_{i}}\right]+\sum_{i}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)$
$=\mathcal{I}(\mathcal{M};\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q}),$
$=\mathcal{H}(\mathcal{M})-H(\mathcal{M}|\Omega)+D_{\Delta}(\mathbf{p}||\mathbf{q})$
@@
$\Phi(\mathcal{M};\Omega)\geq\mathcal{I}(\mathcal{M};\Omega),$ with equality if and only if $\mathbf{p}=\mathbf{q}$ [^6] This corollary resolves a confusion in [3] by showing that pragmatic information and mutual information are equal only in this special case..
When $\mathbf{p}=\mathbf{q}$, the information needed to specify outcome $\omega_{i}\in\Omega$, i.e. the amount of information used in making a presumably useful decision, is the total amount of information transmitted by receipt of $m\in\mathcal{M}$, reduced by $H(\mathcal{M}|\Omega)$, the amount of information that is unneeded to specify the given outcome, and thus useless for making the decision. This difference is similar in spirit to free energy, the difference between the total energy and the internal energy, i.e. the information that is not available to do useful work. We can therefore characterize pragmatic information as “free information”; in fact, it is a generalization of the free information defined in [1].
In the more general case in which $\mathbf{p}\neq\mathbf{q}$, the term $D_{\Delta}(\mathbf{p}||\mathbf{q})$ is needed to update the biases that $\Delta$ had prior to the receipt of $m\in\mathcal{M}$. In this case, $\Delta$ is provided with two kinds of information; in addition to $\mathcal{H}(\mathcal{M})$, $\Delta$ is also provided with the information of the update. The characterization of pragmatic information as free information, i.e. the difference between total information and “irrelevant information”, still makes sense.